Solve by the substitution method.
3x + 4y = 20
-6x + y = 32
Will someone please help me with this one.

Answer by scott8148(6628)

adding 6x to the 2nd equation __ y=32+6x

substituting into the 1st equation __ 3x+4(32+6x)=20

distributing __ 3x+128+24x=20 __ subtracting 128 __ 27x=-108 __ dividing by 27 __ x=-4

substituting __ y=32+[6(-4)] __ y=32-24 __ y=8